An astronaut with a mass of 75kg slides down a slope that is 3,5m high and 25cm from the ground. Assume that the frictional force is negligible. What is the velocity of the astronaut at the end of the slope?


1
Expert's answer
2021-07-19T09:51:44-0400
"v=\sqrt{2gh}\\v=\sqrt{2(9.8)(3.5)}=8.3\frac{m}{s}"